OpenStudy (loser66):

A sky diver weighing 180lb falls vertically downward from an altitude of 5000ft, and opens the parachute after 10sec of free fall. Assume that the force of air resistance is 0.75|v| when the parachute is closed and 12|v| when the parachute is open, where the velocity v is measured in ft/sec. a) Find the speed of the sky diver when the parachute opens b) Find the distance fallen before the parachute opens c)What is the limiting velocity V_l after the parachute opens? d) determine how long the sky driver is in the air after the parachute opens Please, help

OpenStudy (loser66):

It's not physics; It's math, differential equations.

OpenStudy (mertsj):

Do you realize that I have not done DE since 1961?

OpenStudy (mertsj):

@satellite73

OpenStudy (loser66):

@rperez36 this is from your class, too. help me back, friend

OpenStudy (anonymous):

yeah I got you let me draw it out

OpenStudy (loser66):

the equation is \[m\frac {dV}{dt}= g - 0.75V\\\frac{dV}{dt}+\frac{0.75}{m}V=g\]

OpenStudy (anonymous):